### v2

# PiKa: Expert-Level Synthetic Datasets for Post-Training Alignment from Scratch

Abstract
High-quality instruction data is critical for LLM alignment, yet existing open-source datasets often lack efficiency, requiring hundreds of thousands of examples to approach proprietary performance. In this work, we find that beyond the widely recognized importance of prompt–response quality, prompt difficulty itself plays a critical role in driving alignment gains. Motivated by this observation, we introduce PiKa, a data-efficient family of expert-level alignment datasets that concentrates supervision on high-difficulty instructions. The PiKa-SFT dataset contains only 30k examples—an order of magnitude fewer than state-of-the-art open datasets like Magpie-Pro. Despite its small size, fine-tuning Llama-3-8B-Base on PiKa-SFT even outperforms the official Llama-3-8B-Instruct model—trained on over 10M proprietary examples—on widely used benchmarks such as AlpacaEval 2.0 and Arena-Hard. We also validate the generalizability of PiKa across the Qwen2.5 series (0.5B–7B), consistently surpassing their official instruction-tuned counterparts. Additionally, we provide 30k high-quality preference optimization examples to further enhance alignment. Our results demonstrate that promising alignment is achievable with significantly reduced data, democratizing access for resource-constrained research. Our code and data will be available at https://github.com/SJY8460/PiKa.
PiKa: Expert-Level Synthetic Datasets for Post-Training Alignment from Scratch
Shangjian Yin1,2   Shining Liang2   Wenbiao Ding2   Yuli Qian2 Zhouxing Shi1   Hongzhi Li2   Yutao Xie2 1University of California, Riverside   2Microsoft AI
1 Introduction
Large language models (LLMs) have rapidly become the foundation of modern AI systems, achieving strong performance across a broad spectrum of instruction-following tasks (Achiam et al., 2023; Meta, 2024; Qwen et al., 2025). A key factor behind this success lies in instruction fine-tuning, which enables models to generalize beyond their pretraining distribution and handle novel user queries effectively. The quality of this instruction tuning, however, is highly dependent on the availability of reliable alignment datasets. Despite the release of powerful open-source models, the datasets used to align models remain largely proprietary. This lack of transparency poses a major obstacle to reproducibility and hinders progress in open research on improving LLM alignment (Xu et al., 2025).
To overcome the difficulty of building high-quality datasets, two major directions have been explored. The first relies on human experts to author and curate instruction–response pairs (Databricks, 2023; Köpf et al., 2023; Zhao et al., 2024; Zheng et al., 2024, 2023), a process that is both time-consuming and labor-intensive (Liu et al., 2024). The second line of work leverages LLMs themselves to automatically generate synthetic data (Ding et al., 2023; Yin et al., 2023; Li et al., 2024a; Sun et al., 2023; Taori et al., 2023; Wang et al., 2023, 2024; Xu et al., 2023a, b; Li et al., 2023a). More recently, efforts have shifted toward increasing the diversity of synthetic datasets and scaling them up to massive collections of conversational examples that aim to approximate human interaction. However, despite such progress, models trained from scratch on these large-scale synthetic corpora often acquire only limited instruction-following ability, and the training process itself requires significant computational resources, making it less accessible for much of the research community.
Do we really need such massive datasets to enhance the instruction-following ability of base models? Intuitively, training on large collections of diverse but overly simple data pairs provides only limited informational gain, as models tend to mimic surface-level patterns rather than generalize to more complex scenarios. Even when prompts and responses are “high-quality”, prompt difficulty largely determines how much depth of knowledge the model utilizes during optimization. In other words, datasets containing more challenging instructions of comparable quality require richer reasoning and more detailed solutions, which can improve performance not only on difficult tasks but also transfer effectively to simpler ones. Thus, the utility of an instruction dataset is not merely a matter of scale, but of how well its prompts capture the complexity needed to guide robust instruction following.
Figure 1: Overview of the PiKa pipeline for synthesizing expert-level alignment data.
Building on these insights, we construct an expert-level synthetic instruction dataset named PiKa, as illustrated in Figure 1. Our approach is inspired by persona-based generation methods (Ge et al., 2025), which prompt LLMs to role-play as specific personas to generate diverse instructions. To further enhance data quality, we rigorously sample and filter persona prompts, removing harmful or policy-violating cases and retaining only complex personas for instruction generation. Unlike many existing datasets, PiKa is fully synthesized through an automated generation pipeline, enabling scalable and consistent data construction. All generated instruction–response pairs are subsequently filtered and re-scored to prioritize difficulty and solution quality, resulting in a dataset dominated by challenging, expert-level prompts paired with detailed solutions. We provide comprehensive statistics, highlight the advantages of PiKa, and present an in-depth analysis in Section 3, enabling practitioners to flexibly select and filter data for fine-tuning according to their specific needs.
To evaluate the effectiveness of PiKa, we compare it with widely used public instruction datasets (e.g., ShareGPT (Chiang et al., 2023), WildChat (Zhao et al., 2024), UltraChat (Ding et al., 2023), OpenHermes (Teknium, 2023b, a), Tulu V2 Mix (Ivison et al., 2023), and Magpie-Air/Pro (Xu et al., 2025)) by conducting supervised fine-tuning (SFT) of the Llama-3-8B-Base model on each dataset and evaluating the resulting models on standard alignment benchmarks such as AlpacaEval 2 (Li et al., 2023b) and Arena-Hard (Li et al., 2024b), following recent Magpie studies. Our results show that models fine-tuned on PiKa consistently achieve superior performance, even surpassing those trained with both SFT and direct preference optimization (DPO) (Rafailov et al., 2023) on UltraFeedback (Cui et al., 2023). Remarkably, PiKa is more than an order of magnitude smaller than Magpie-Pro (30K vs. 300K examples for SFT). Despite this significant difference in dataset size, PiKa achieves superior performance across benchmarks. For the first time, our approach enables open-source aligned models to significantly outperform the official Llama-3-8B-Instruct on both AlpacaEval 2.0 and Arena-Hard. This is particularly remarkable given that Llama-3-8B-Instruct was trained with over 10M proprietary examples and subsequent preference optimization. Building on this, we extend our study to the Qwen2.5 (0.5B–7B), where models trained on PiKa-SFT consistently surpass their officially tuned counterparts.
Furthermore, PiKa not only excels under SFT alone compared to prior public datasets, but also delivers the best results when combined with preference optimization methods such as DPO. By extending PiKa to generate high-quality preference optimization data, PiKa-aligned Llama-3 models achieve substantial improvements and outperform models trained with both the Magpie and UltraFeedback series, with particularly strong gains on the more challenging Arena-Hard benchmark (43.70 vs. 33.30 for Magpie-Pro and 25.30 for UltraFeedback respectively). These findings highlight the exceptional quality and efficiency of PiKa, demonstrating that synthetic data constructed by prioritizing prompt difficulty and expert-level supervision can outperform extensively optimized proprietary datasets in a far more data-efficient manner.
2 PiKa: Expert-Level Synthetic Datasets
Our primary contribution is the construction of PiKa, a family of expert-level synthetic datasets for data-efficient alignment. This section introduces the design principles and generation framework underlying the PiKa datasets.
Step 1: Expert-Level Instruction Generation. We sample a diverse set of personas $\{\pi_{i}\}_{i=1}^{N}$ from PersonaHub (Ge et al., 2025), covering multiple domains such as biology, engineering, medicine, and law. Each persona $\pi_{i}$ is then provided to GPT-4o, which auto-regressively generates a corresponding knowledge-intensive instruction. Based on our preliminary experiments, we find that more domain-specific and sophisticated personas tend to generate more challenging prompts:
$I_{i}=\text{LLM}(\pi_{i}),\quad i=1,\dots,N.$ (1)
To ensure diversity, each persona is used only once. Generated instructions are further filtered through a quality control sampling $Q(I_{i})\in\{0,1\}$, retaining only those with $Q(I_{i})=1$, i.e., instructions that are challenging, safe, and informative.
Step 2: Multi-Path Response Generation. For each validated instruction $I_{i}$, the LLM produces $k$ candidate responses under mild stochasticity with temperature $T<1$:
$R_{i}=\{r_{i1},r_{i2},\dots,r_{ik}\}=\text{LLM}(I_{i};T).$ (2)
This yields diverse data pairs $(I_{i},r_{ij})$ that differ in reasoning depth, style, and completeness.
Step 3: Reward-Model-Guided Selection. For each instruction $I_{i}$ ($i=1,\dots,N$), a reward model $\mathcal{R}(\cdot)$ assigns a score to each candidate responses. In our default setting, we employ Skywork-Reward-V2-Llama-3.1-8B (Liu et al., 2025), which ranks first on the RewardBench leaderboard [^1] https://huggingface.co/spaces/allenai/reward-bench, as of 2026/4/8. The reward score for each candidate response $r_{ij}$ of instruction $I_{i}$ is computed as:
$s_{ij}=\mathcal{R}(I_{i},r_{ij}),\quad j=1,\dots,k,$ (3)
where $k$ is the number of candidate responses per instruction.
For SFT, we retain the pair $(I_{i},r_{ij^{*}})$ with
$j^{*}=\arg\max_{j}s_{ij}.$ (4)
For preference optimization (e.g., DPO), we construct triples $\big(I_{i},r_{ij^{+}},r_{ij^{-}}\big)$ by selecting the highest- and lowest-scoring responses:
$j^{+}=\arg\max_{j}s_{ij},\quad j^{-}=\arg\min_{j}s_{ij}.$ (5)
These preference pairs provide supervision for optimizing models toward expert-level alignment.
3 Dataset Analysis
We conduct a comprehensive analysis of the proposed PiKa dataset, covering instruction similarity, length statistics, difficulty, quality, feasibility, and construction cost. This study demonstrates that PiKa delivers more challenging, diverse, and practically useful supervision compared to existing datasets such as Magpie-Pro.
Figure 2: (a) Minimum input embedding distances distribution. (b) Instruction lengths distribution. (c) Response lengths Distribution.
3.1 Instruction Similarity
To evaluate dataset coverage, we follow prior work (Zhao et al., 2024; Xu et al., 2025) and compute the minimum neighbor distance (MND) in the embedding space. Instructions are encoded with all-mpnet-base-v2 , and pairwise distances are calculated with FAISS (Douze et al., 2024). The MND for an instruction $I_{i}$ is defined as:
$\text{MND}(I_{i})=\min_{j\neq i}\;\|e(I_{i})-e(I_{j})\|_{2},$ (6)
where $e(I)$ denotes the embedding of instruction II. Larger MND indicates lower redundancy and higher diversity. As shown in Figure 2(a), PiKa achieves a mean MND of 0.497 compared to Magpie-Pro’s 0.598, indicating comparable diversity while effectively avoiding repetitive prompts that plague many existing datasets.
3.2 Length Analysis
We further analyze the tokenized lengths of instructions and responses. Figure 2(b, c) shows that PiKa instructions average around 424 characters, and responses exceed 5,305 characters (around 1300 tokens) on average—substantially longer than existing datasets—indicating that PiKa provides more detailed and knowledge-intensive supervision.
3.3 Dataset Assessment
Beyond coverage and length, we adopt GPT-4o as an expert evaluator to assess difficulty, feasibility, and quality. Each attribute is scored on a 1–10 scale using dedicated prompts (see Appendix B). Formally, we define $s_{d}:\mathcal{X}\!\to\![1,10]$, $s_{f}:\mathcal{X}\!\to\![1,10]$, and $s_{q}:\mathcal{X}\!\times\!\mathcal{Y}\!\to\![1,10]$, which denote the difficulty of instruction $x$, the feasibility of instruction $x$, and the quality of an instruction–response pair $(x,y)$, respectively.
Figure 3: LLM–based evaluation of PiKa’s instruction difficulty, feasibility and instruction-response quality compared with Magpie-Pro. We also provide results from multiple judges in Section E.7.
3.3.1 Difficulty Evaluation
We evaluate instruction difficulty based on task complexity, required domain knowledge, cognitive load, and technical sophistication. The framework categorizes tasks into elementary knowledge (scores 1–2), specialized expertise (3–6), expert-level understanding (7–8), and cutting-edge research knowledge (9–10). As shown in Figure 3(a) , PiKa instructions exhibit significantly higher difficulty, with a mean score of 7.39 compared to Magpie-Pro’s 2.65. The distribution further reveals that while Magpie-Pro concentrates heavily in the elementary to intermediate range (scores 1–4), PiKa covers a broader spectrum, with a substantial proportion requiring expert-level reasoning and advanced technical understanding. This demonstrates that PiKa effectively captures more sophisticated tasks, and Figure 6 provides illustrative comparisons.
3.3.2 Feasibility Evaluation
Feasibility assessment examines whether instructions represent realistic, executable tasks with practical applicability. The evaluation considers task feasibility, contextual reasonableness, logical consistency, and real world relevance. Scores range from completely unrealistic tasks (1 to 2) to highly practical and useful instructions (9 to 10). As demonstrated in Figure 3(b), PiKa maintains strong feasibility with a mean score of 8.98 compared to Magpie-Pro’s 8.77. Notably, both datasets show high concentration in the upper feasibility ranges (scores 8 to 10), but PiKa demonstrates a slightly more consistent distribution in the high feasibility region. This ensures that despite the increased difficulty, PiKa instructions remain grounded in practical, real world applications that users would genuinely encounter.
3.3.3 Quality Evaluation
Response quality is evaluated based on accuracy, completeness, clarity, instruction response alignment, educational value, and real world applicability. The scoring ranges from poor quality with incorrect information (1 to 2) to excellent quality with comprehensive and insightful responses (9 to 10). Figure 3(c) shows that PiKa achieves superior response quality with a mean score of 9.57 compared to Magpie-Pro’s 9.01. Both datasets demonstrate high quality responses concentrated in the 8 to 10 range, but PiKa shows a more pronounced peak at score 10, indicating a higher proportion of exceptional responses. The quality improvement, combined with significantly longer response lengths, suggests that PiKa responses not only meet accuracy standards but also provide more thorough, comprehensive, and educationally valuable content for alignment training.
Table 1: Performance comparison of instruction-tuned models based on Llama-3-8B-Base using PiKa-generated versus baseline datasets. PiKa achieves superior performance while requiring $10\times$ less training data than state-of-the-art Magpie methods. Standard deviations (SD) reported are computed across both LC and WR metrics and #Convs indicates the number examples used for training the base model.
Alignment Setup (Base LLM = Llama-3-8B-Base) #Convs AlpacaEval 2 Arena-Hard
GPT-4-Turbo (1106) Llama-3-8B-Instruct GPT-4 (0314)
LC (%) WR (%) SD LC (%) WR (%) SD WR(%)
Llama-3-8B-Instruct (Official) >>10M 28.36 27.93 1.48 50.0 50.0 - 24.5
Self-Instruct (Llama-3) (Wang et al., 2023) 100K 8.86 4.16 0.66 24.48 11.97 1.09 3.3
ShareGPT (Chiang et al., 2023) 112K 6.98 4.00 0.64 26.05 15.98 1.22 6.9
Ultrachat (Ding et al., 2023) 208k 6.70 3.48 0.60 24.14 13.37 1.14 3.6
OpenHermes 1 (Teknium, 2023b) 243K 8.69 4.98 0.70 26.81 16.98 1.25 5.3
Tulu V2 Mix (Ivison et al., 2023) 326k 10.95 6.43 0.81 24.84 15.39 1.20 6.3
WildChat (Zhao et al., 2024) 652K 14.75 8.43 0.91 33.99 23.77 1.42 11.7
OpenHermes 2.5 (Teknium, 2023a) 1M 12.40 7.80 0.87 37.14 27.86 1.48 7.7
Magpie-Air-300K-Filtered (Xu et al., 2025) 300K 25.24 27.33 1.47 43.79 47.35 1.69 20.7
Magpie-Pro-300K-Filtered (Xu et al., 2025) 300K 24.06 28.60 1.51 41.28 46.80 1.68 23.9
PiKa (Ours) 30K 32.82 30.56 1.54 52.42 50.30 1.68 33.5
Table 2: Performance comparison of models instruction-tuned on Qwen2.5 base models using the PiKa 30K dataset versus official instruction-tuned models. PiKa consistently outperforms official models across all model sizes.
Alignment Setup AlpacaEval 2
GPT-4-Turbo (1106) Official Aligned Model as Ref.
LC (%) WR (%) SD LC (%) WR (%) SD
Qwen2.5-0.5B Qwen2.5-0.5B-Instruct 1.30 0.93 0.32 50 50 -
Base Model + PiKa 1.52 2.15 0.49 55.29 55.65 1.66
Qwen2.5-1.5B Qwen2.5-1.5B-Instruct 3.77 2.21 0.48 50 50 -
Base Model + PiKa 9.64 11.07 1.04 70.68 74.35 1.47
Qwen2.5-3B Qwen2.5-3B-Instruct 6.82 5.16 0.72 50 50 -
Base Model + PiKa 14.78 15.83 1.23 63.25 66.97 1.56
Qwen2.5-7B Qwen2.5-7B-Instruct 29.03 25.99 1.47 50 50 -
Base Model + PiKa 32.55 33.24 1.59 58.87 60.23 1.65
4 Performance Analysis
4.1 Experimental Setup
Baselines for Supervised Fine-Tuning and Preference Optimization. We compare PiKa with nine state-of-the-art open-source instruction datasets: Self-Instruct (Wang et al., 2023), ShareGPT (Chiang et al., 2023), WildChat (Zhao et al., 2024), UltraChat (Ding et al., 2023), OpenHermes 1 (Teknium, 2023b), OpenHermes 2.5 (Teknium, 2023a), Tulu V2 Mix (Ivison et al., 2023), Magpie-Air (Xu et al., 2025), and Magpie-Pro (Xu et al., 2025). ShareGPT and WildChat represent human-authored datasets containing 112K and 652K high-quality multi-turn conversations between humans and ChatGPT, respectively. UltraChat and the Magpie family are representative open-source synthetic datasets. For preference optimization, we compare models aligned using PiKa with direct preference optimization (DPO) (Rafailov et al., 2023) baselines, specifically comparing against two state-of-the-art open-source training datasets: UltraFeedback and Magpie-Pro.
Model Alignment Details. We conduct experiments on the Llama-3 and Qwen2.5 base models. For supervised fine-tuning, we follow Touvron et al. (2023) and employ a cosine learning rate schedule with an initial learning rate of $2\times 10^{-5}$. The maximum sequence length is set to 8192 tokens. For DPO training, we use a cosine learning rate of $5\times 10^{-7}$. We adhere to the official instruction templates of each respective model architecture.
Evaluation Benchmarks. We evaluate aligned model performance using two widely adopted instruction-following benchmarks: AlpacaEval 2 (Li et al., 2023b) and Arena-Hard (Li et al., 2024b). AlpacaEval 2 comprises 805 representative instructions selected from real user interactions, providing a comprehensive assessment of practical instruction-following capabilities. Arena-Hard represents an enhanced version of MT-Bench (Zheng et al., 2023), containing 500 challenging user queries designed to test advanced reasoning and problem-solving abilities. Both benchmarks employ GPT-based evaluators to assess responses generated by the model under evaluation against established baseline models. Specifically, we use GPT-4-Turbo (1106) and Llama-3-8B-Instruct as baselines for AlpacaEval 2 and Arena-Hard uses GPT-4 (0314) as its baseline. In addition, we primarily employ the latest GPT-4o as the LLM-as-judge to provide contemporary and robust evaluation standards. We also include evaluations using GPT-5 and GPT-4.1-mini, as shown in Table 4.
Metrics. We adopt two complementary metrics to measure instruction-following capabilities of fine-tuned models. The first metric is the win rate (WR), which calculates the fraction of responses favored by the GPT evaluator over the baseline model. This metric is applied across both AlpacaEval 2 and Arena-Hard benchmarks. The second metric is the length-controlled win rate (LC) (Dubois et al., 2024), a debiased variant of WR that accounts for response length differences. The GPT evaluator considers the lengths of responses generated by both the baseline model and the model under evaluation when computing LC, thereby reducing the potential bias introduced by response length variations. This metric is specifically applied to the AlpacaEval 2 benchmark (Li et al., 2023b).
Table 3: Ablation study on prompt difficulty. We compare the Magpie-Pro baseline with different PiKa variants. We also provide an ablation on the generator choice for PiKa in Section E.9.
AlpacaEval 2 Dataset Difficulty LC (%) WR (%) Magpie-Pro 2.65 15.42 16.89 PiKa-Series (10k Subset)      w/o Persona-Guide 3.11 13.84 15.53      Low-Diff 2.91 21.86 14.95      Mid-Diff 3.64 24.36 17.84      Expert (Default) 7.39 31.01 30.32
Table 4: Additional evaluation results on AlpacaEval 2 judged by GPT-5 and GPT-4.1-mini. Note: LC is the most important metric.
AlpacaEval 2 Judger Model LC (%) WR (%) GPT-5 Official-Instruct 15.95 16.56 Magpie-Pro 9.51 14.55 PiKa 19.28 19.00 GPT-4.1-mini Official-Instruct 25.44 25.53 Magpie-Pro 20.17 28.54 PiKa 29.02 27.55
Table 5: Performance comparison on additional downstream objective tasks from the Open LLM Leaderboard. The goal of this evaluation is to assess whether alignment with PiKa preserves performance on objective tasks, rather than to optimize for further gains. All models are supervised fine-tuned on Llama-3-8B-Base. Numbers in parentheses indicate the number of few-shot examples.
Alignment Setup MMLU (5) ARC (25) HellaSwag (10) TruthfulQA (0) WinoGrande (5) GSM8K (5) Average
Llama-3-8B-Instruct 67.82 61.52 78.67 52.47 72.14 71.72 67.39
ShareGPT 66.03 58.45 81.50 52.34 74.03 48.67 63.50
OpenHermes 1 65.42 62.29 82.15 50.85 75.61 47.16 63.58
OpenHermes 2.5 65.70 61.86 82.53 51.35 76.09 67.02 67.09
Tulu V2 Mix 66.34 59.22 82.80 47.99 76.16 58.07 65.10
WildChat 65.95 59.22 81.39 53.18 75.30 48.75 63.97
UltraChat 65.23 62.12 81.68 52.76 75.53 50.57 64.65
Magpie-Air-300K-Filtered 64.45 61.01 79.90 53.48 72.38 52.24 63.58
Magpie-Pro-300K-Filtered 64.25 60.41 80.52 52.46 73.32 47.92 63.15
PiKa 62.85 59.98 80.02 52.48 73.01 52.84 63.53
4.2 Experimental Results
[…]
PiKa surpasses official aligned models with significantly fewer data. When compared against the official Llama-3-8B-Instruct model (which underwent extensive SFT and DPO training on over 10M conversations), PiKa demonstrates exceptional efficiency. Our model achieves 52.42% LC and 50.30% WR when using Llama-3-8B-Instruct as the reference, indicating a clear preference over the officially aligned version. This represents a significant improvement over the 50% baseline, demonstrating the exceptional quality of PiKa-generated data. The Arena-Hard results further corroborate this finding, with PiKa achieving 33.5% WR compared to the official model’s 24.5%.
[…]
Figure 4: Comparison across three dataset families: UltraFeedback, Magpie-Pro, and PiKa. .
(a) AlpacaEval 2.0 LC scaling
(b) AlpacaEval 2.0 WR scaling
(c) Arena-Hard WR scaling
Figure 5: Scaling analysis of dataset size from 10K to 30K comparing PiKa versus Magpie-Pro. PiKa demonstrates consistent superiority across all data scales and maintains better performance curves, with optimal performance achieved at 30K examples on Arena-Hard.
Expert-level prompt difficulty is a key driver of performance gains. To verify the effectiveness of the difficulty-driven construction strategy, we investigate the impact of prompt difficulty on the model’s generative chat ability. We trained models on 10k subsets across different categories. As shown in Table 3, the w/o Persona-Guide variant fails to surpass the baseline, indicating that difficulty alone is insufficient without persona constraints. Crucially, we observe a positive correlation between prompt difficulty and chat performance. As the difficulty score rises from 2.91 (Low-Diff) to 3.64 (Mid-Diff) and finally to 7.39 (Expert), the generative chat ability improves significantly. Specifically, the Expert subset boosts LC from 21.86% to 31.01% and doubles the WR from 14.95% to 30.32%, suggesting that prompt difficulty is a key driver for superior chat capabilities.
Preference optimization further amplifies performance gains. To evaluate PiKa’s effectiveness in preference optimization, we compare it against UltraFeedback and Magpie-Pro across data efficiency and benchmark performance (Figure 4). As shown in Figure 4(a), PiKa is highly data-efficient, requiring only 30K examples for SFT and 30K for DPO. In contrast, UltraFeedback uses 208K and 64K, while Magpie-Pro uses 300K and 60K, respectively. Despite this substantial reduction in training data, PiKa achieves superior results on both evaluation benchmarks. On AlpacaEval 2.0, PiKa attains a 41.29% win rate, significantly surpassing Magpie-Pro (36.24%) and UltraFeedback (18.40%). Similarly, on Arena-Hard, PiKa achieves 43.70%, outperforming Magpie-Pro (33.30%) and UltraFeedback (25.30%). These results highlight PiKa’s ability to generate high-quality preference data, enabling more effective alignment with human preferences even under constrained data budgets.
Scaling Analysis. Figure 5 presents a scaling analysis comparing PiKa and Magpie-Pro across dataset sizes from 10K to 30K. Across all metrics and scales, PiKa consistently outperforms Magpie-Pro. Notably, PiKa with only 10K examples achieves performance surpassing, Magpie-Pro with 30K examples. The scaling curves further show that PiKa reaches its peak performance at 30K examples on Arena-Hard. We therefore adopt the 30K dataset as the default setting due to its highest Arena-Hard score. Nevertheless, future work could focus more on data efficiency, as PiKa with 10K examples already demonstrates competitive performance.
Performance on Additional Benchmarks. Table 5 reports performance across diverse tasks from the Hugging Face Open LLM Leaderboard (Beeching et al., 2023), including MMLU (Hendrycks et al., 2020), ARC Challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), TruthfulQA (Lin et al., 2021), WinoGrande (Levesque et al., 2012), and GSM8K (Cobbe et al., 2021). PiKa achieves competitive performance (63.53% average) comparable to other datasets, demonstrating that the quality improvements in instruction-following do not come at the expense of general capabilities. The slight performance difference on mathematical reasoning tasks (eg., GSM8K) can be attributed to the limited proportion of mathematical instructions in PiKa’s current composition. Future iterations can incorporate domain-specific mathematical and coding datasets to enhance performance on reasoning-intensive tasks while maintaining PiKa’s strengths in general instruction-following.
5 Related Work
LLM Alignment. A central goal in large language model (LLM) research is aligning model behavior with human values and intentions. Two major approaches are instruction tuning and preference tuning. Instruction tuning (Wei et al., 2022) fine-tunes LLMs on datasets consisting of user instructions and target responses, which can be single- or multi-turn. The effectiveness of this method is closely tied to the quality and diversity of the instruction data (Taori et al., 2023; Wang et al., 2023; Zhou et al., 2023). Preference tuning, on the other hand, builds upon instruction tuning by further refining model outputs using either reinforcement learning from human feedback (RLHF) (Bai et al., 2022) or direct preference optimization methods (Azar et al., 2024; Ethayarajh et al., 2024; Hong et al., 2024; Rafailov et al., 2023). These approaches leverage datasets containing preference comparisons, enabling models to better capture nuanced human judgments.
Persona Roleplay. Persona-based roleplay offers a powerful mechanism for exploring the knowledge and perspectives embedded in LLMs. In this paradigm, the model’s world knowledge is effectively “compressed” into distributed representations, which can then be “decompressed” through the adoption of diverse personas to generate contextually grounded text (delétang2024languagemodelingcompression; Ge et al., 2024). Recent work by Ge et al. (2025) introduced the Persona Hub, an automatically constructed repository derived from large-scale web data. This resource allows researchers to tap into a wide variety of perspectives within LLMs, facilitating the large-scale creation of synthetic data without reliance on a small fixed seed corpus. Compared to traditional prompt-based synthesis, persona-driven approaches can be seamlessly integrated into arbitrary prompts, leveraging the strong role-playing capacity of LLMs to generate more diverse and versatile data. In our work, persona signals are randomly sampled from the Persona Hub to construct expert-level prompts.
Dataset Construction. Methods for building alignment datasets generally fall into two categories: human-in-the-loop collection and synthetic instruction generation. The first line of work gathers datasets through human interactions with LLMs (Databricks, 2023; Zhao et al., 2024; Zheng et al., 2024, 2023; Köpf et al., 2023). While such data is often high quality, it is expensive to scale, time-consuming to curate, and may contain toxic or biased content (Zhao et al., 2024). The second line of work leverages LLMs to generate synthetic instructions from a small set of human-annotated seeds, which are expanded using prompting strategies (Wang et al., 2023; Taori et al., 2023; Xu et al., 2023a, b; Wang et al., 2024; Sun et al., 2023). However, these methods often suffer from limited diversity, as generated instructions tend to resemble the original seeds (Li et al., 2024a). To address this, some approaches attempt to broaden coverage by leveraging summaries of world knowledge to drive generation (Ding et al., 2023; Li et al., 2024a). Our proposed PiKa dataset belongs to the synthetic category, but differs in that we employ persona-driven prompts sampled from expert-level profiles and selectively retain high-difficulty instances to better capture preference signals.
6 Conclusion
We present PiKa, a family of expert-level alignment datasets for LLMs. By leveraging persona-driven instruction generation and automatic filtering, we curated a compact yet diverse set of SFT and preference data that improves alignment quality. Fine-tuning experiments on Llama-3-8B show that models trained on PiKa-SFT outperform those trained on much larger open-source datasets and even surpass the official Llama-3-8B-Instruct model, trained on over 10M proprietary examples. We further validate PiKa’s effectiveness across different model families, consistently outperforming their official instruction-tuned counterparts. These results highlight PiKa’s potential to make alignment research more data-efficient for the open-source community.
7 Limitation
The current version of PiKa does not include math or code reasoning data. We plan to iterate on this framework to incorporate such data in future versions. Additionally, the PiKa dataset has significant potential for size reduction. Future work can focus on more sophisticated data selection strategies to further reduce the training dataset, as we observe that 10k PiKa examples can achieve comparable performance to 30k examples during training.
Appendix A Ethics Statement
We carefully synthesize data using GPT-4o and filter out unsafe, harmful, or sensitive data pairs during the initial generation process. To the best of our knowledge, the resulting dataset does not contain any content that could cause harm or violate ethical standards.
Appendix B Evaluation Prompt
B.1 Difficulty evaluation prompt
We employ the following system prompt to evaluate the difficulty level of instructions:
System Prompt: You are an expert evaluator for assessing the difficulty level of instructions/questions. Please evaluate the given instruction on a scale of 1-10 based on: Complexity of the task or question Level of domain knowledge required Cognitive load needed to understand and process Technical depth and sophistication Scoring guidelines: 1-2: Very basic, elementary level knowledge 3-4: Intermediate, requires some specialized knowledge 5-6: Advanced, requires solid domain expertise 7-8: Expert level, requires deep specialized knowledge 9-10: Cutting-edge, requires highly specialized or research-level expertise Please respond with ONLY a single number (1-10) representing the difficulty score.
B.2 Realism evaluation prompt
We employ the following system prompt to evaluate the realism and feasibility of instructions:
System Prompt: You are an expert evaluator for assessing the realism and feasibility of instructions/prompts. Please evaluate the given instruction on a scale of 1-10 based on: Feasibility: Can this task actually be completed in the real world? Practicality: Does this request make sense in real-world scenarios? Realistic context: Are the assumptions and requirements reasonable? Real-world applicability: Would someone actually need this in practice? Logical consistency: Are there any contradictions or impossible requirements? Scoring guidelines: 1-2: Completely unrealistic, impossible to execute, or nonsensical 3-4: Mostly unrealistic, major feasibility issues or impractical requirements 5-6: Somewhat realistic but has questionable assumptions or limited applicability 7-8: Quite realistic and feasible, minor issues but generally practical 9-10: Highly realistic, completely feasible, and practically useful Please respond with ONLY a single number (1-10) representing the realism score.
B.3 Quality evaluation prompt
We employ the following system prompt to evaluate the quality of instruction-response pairs:
System Prompt: You are an expert evaluator for assessing the quality of instruction-response pairs. Please evaluate the given instruction-response pair on a scale of 1-10 based on: Accuracy and correctness of the response Completeness and thoroughness of the answer Clarity and coherence of explanation Alignment between instruction and response Educational value and helpfulness Reflection of human preferences and real-world applicability Scoring guidelines: 1-2: Poor quality, incorrect, unhelpful, or completely misaligned 3-4: Below average, partially correct but lacks clarity or completeness 5-6: Average quality, generally correct but could be improved 7-8: Good quality, accurate, clear, and helpful 9-10: Excellent quality, comprehensive, insightful, and highly valuable Please respond with ONLY a single number (1-10) representing the quality score.
Appendix C Examples from PiKa Dataset
To illustrate the application of our evaluation framework, we present two representative examples from the PiKa dataset with their corresponding instruction-response pairs.
C.1 Example 1: Complex Scientific Analysis
Instruction:
Provide an in-depth analysis of non-lethal whale research methods used in the Southern Ocean, including their effectiveness compared to lethal research practices. Discuss the ethical, ecological, and scientific advantages of these methods, and provide examples of successful non-lethal whale studies. Additionally, examine Japan’s whaling practices from a conservation perspective and explain how international cooperation, particularly through the International Whaling Commission, can promote sustainable whale management.
Response:
Non-Lethal Whale Research Methods in the Southern Ocean: An In-Depth Analysis
Whale research in the Southern Ocean has evolved significantly, with non-lethal methods becoming increasingly effective alternatives to lethal whaling practices. Non-lethal techniques provide valuable scientific data while preserving whale populations and maintaining ecological balance. These methods are widely recognized for their ethical, ecological, and scientific advantages over lethal research practices.
Non-Lethal Research Methods
Several non-lethal approaches have been successfully implemented in whale research, demonstrating their effectiveness in studying whale populations, behavior, and health:
1. Biopsy Sampling
  • Small tissue samples are obtained using a specialized dart system.
  • Provides data on genetics, hormone levels, pollutants, and diet.
  • Less invasive and does not harm the whale.
2. Photo Identification
  • Researchers track individual whales using unique markings and patterns on their bodies.
  • Allows long-term monitoring of population dynamics and migration patterns.
  • Successfully used to estimate population sizes and reproductive rates.
3. Acoustic Monitoring
  • Hydrophones record whale vocalizations to study communication and population distribution.
  • Helps assess the impact of noise pollution on whale behavior.
  • Provides real-time data on species presence and movement.
4. Satellite Tagging
  • Non-invasive tags are attached to whales to track migration patterns and feeding behavior.
  • Offers insights into habitat use and climate change impacts on whale movements.
  • Long-term tracking improves conservation strategies.
5. Environmental DNA (eDNA) Analysis
  • Water samples are analyzed for genetic material shed by whales.
  • Enables species identification and population assessments without direct contact.
  • Emerging as a highly effective and non-invasive research tool.
Effectiveness Compared to Lethal Research Practices
Non-lethal methods provide comprehensive scientific data without harming whale populations. Lethal research, historically justified for studying whale age, diet, and reproductive health, has been largely discredited due to advancements in non-lethal techniques. The effectiveness of non-lethal methods is evident in their ability to:
  • Yield high-quality genetic, ecological, and behavioral data.
  • Allow long-term monitoring of individual whales.
  • Avoid disrupting population stability and ecosystem balance.
Ethical, Ecological, and Scientific Advantages
Ethical Benefits
  • Eliminates unnecessary harm and suffering.
  • Aligns with global conservation and animal welfare principles.
  • Enhances public support for whale research and conservation.
Ecological Benefits
  • Preserves whale populations, ensuring their role in marine ecosystems.
  • Prevents disruptions in food chains and biodiversity.
  • Supports sustainable ocean management efforts.
Scientific Benefits
  • Provides more reliable and replicable data over extended periods.
  • Facilitates international collaboration in whale conservation.
  • Advances knowledge of whale behavior and adaptation to environmental changes.
Successful Non-Lethal Whale Studies
Several studies highlight the success of non-lethal research in the Southern Ocean:
  • The Antarctic Humpback Whale Project (AHWP): Uses photo identification and satellite tagging to monitor whale populations.
  • Southern Ocean Research Partnership (SORP): A multinational initiative employing acoustic monitoring and biopsy sampling.
  • Australian Antarctic Division’s Whale Research Program: Uses eDNA analysis and aerial surveys to assess whale distribution.
Japan’s Whaling Practices from a Conservation Perspective
Japan has historically conducted whaling under the guise of scientific research, particularly in the Southern Ocean. Despite the 1986 International Whaling Commission (IWC) moratorium on commercial whaling, Japan has continued whaling operations through loopholes in the IWC framework.
Conservation Concerns
  • Population Decline: Targeting endangered species, such as minke and fin whales, threatens population recovery.
  • Ecosystem Disruption: Whales play a crucial role in nutrient cycling, and their removal affects marine biodiversity.
  • Questionable Scientific Justification: Most data obtained through lethal whaling can be collected using non-lethal methods.
International Cooperation and the Role of the IWC
The International Whaling Commission (IWC) plays a vital role in promoting sustainable whale management through international cooperation. Key strategies include:
1. Strengthening International Regulations
  • Enforcing the moratorium on commercial whaling.
  • Closing loopholes that allow scientific whaling to continue.
  • Implementing stricter penalties for non-compliance.
2. Encouraging Non-Lethal Research
  • Providing funding and support for non-lethal whale studies.
  • Facilitating data-sharing initiatives among member nations.
  • Promoting ethical research practices.
3. Diplomatic and Economic Pressure
  • Encouraging Japan and other whaling nations to transition to whale-watching industries.
  • Leveraging trade agreements to incentivize compliance with conservation efforts.
  • Engaging in diplomatic negotiations to phase out unnecessary whaling.
4. Public Awareness and Advocacy
  • Increasing global awareness of whale conservation issues.
  • Supporting NGOs and scientific communities advocating for non-lethal research.
  • Encouraging responsible ecotourism as an alternative to whaling.
Conclusion
Non-lethal whale research methods in the Southern Ocean have proven to be highly effective, offering ethical, ecological, and scientific advantages over lethal practices. Japan’s continued whaling poses significant conservation challenges, but international cooperation through the IWC and other global initiatives can promote sustainable whale management. By prioritizing non-lethal research and enforcing conservation regulations, the global community can ensure the protection and long-term survival of whale populations.
C.2 Example 2: Medical Information Request
Instruction:
Explain the common causes of eye twitching and provide practical strategies for treatment and prevention, including relaxation techniques, dietary adjustments, and medical interventions. Additionally, discuss rare cases where eye twitching persists despite treatment and when surgery may be considered as a last resort.
Response:
Common Causes of Eye Twitching
Eye twitching, also known as myokymia, is an involuntary spasm of the eyelid muscles. It is usually benign and temporary, but persistent twitching can be bothersome. Common causes include:
  1. 1.
    Fatigue and Lack of Sleep – Insufficient rest can lead to muscle spasms, including twitching in the eyelid.
  2. 2.
    Stress and Anxiety – Emotional strain can trigger involuntary muscle movements.
  3. 3.
    Caffeine and Alcohol Consumption – Excessive intake of stimulants can overexcite the nervous system, leading to twitching.
  4. 4.
    Eye Strain – Prolonged screen time or reading without breaks can fatigue the eye muscles.
  5. 5.
    Dry Eyes – A lack of moisture can irritate the eye and contribute to twitching.
  6. 6.
    Nutritional Deficiencies – Lack of essential nutrients, such as magnesium, potassium, or vitamin B12, can cause muscle spasms.
  7. 7.
    Allergies – Eye irritation from allergens can lead to twitching.
  8. 8.
    Medication Side Effects – Some medications, including those for epilepsy and mental health conditions, may cause eye twitching.
Treatment and Prevention Strategies
1. Relaxation Techniques
  • Stress Management: Engage in activities like meditation, deep breathing, or yoga to reduce stress.
  • Adequate Sleep: Ensure you get 7-9 hours of sleep per night to allow your body to recover.
  • Eye Rest: Follow the 20-20-20 rule—every 20 minutes, look at something 20 feet away for 20 seconds to reduce eye strain.
2. Dietary Adjustments
  • Reduce Caffeine and Alcohol: Cutting back on coffee, tea, and alcohol can help minimize twitching.
  • Increase Magnesium Intake: Magnesium-rich foods (nuts, spinach, bananas) may help prevent muscle spasms.
  • Stay Hydrated: Dehydration can contribute to twitching, so drink plenty of water.
  • Consume Enough Vitamin B12 and Potassium: These nutrients support nerve function and muscle health.
3. Medical Interventions
  • Artificial Tears: If dry eyes are a contributing factor, lubricating eye drops can help.
  • Corrective Lenses: If eye strain is due to vision problems, glasses or contacts may be necessary.
  • Botox Injections: In cases of chronic twitching, botulinum toxin can temporarily relieve symptoms by relaxing the muscles.
  • Medication Adjustments: If twitching is a side effect of a prescribed drug, consult your doctor about alternatives.
Persistent Eye Twitching and Surgical Considerations
In rare cases, eye twitching may persist despite treatment, potentially indicating an underlying neurological disorder such as benign essential blepharospasm or hemifacial spasm. If conservative treatments fail, surgery may be considered:
  • Myectomy: A surgical procedure that removes some of the muscles responsible for twitching.
  • Microvascular Decompression (for Hemifacial Spasm): A neurosurgical procedure that relieves pressure on facial nerves.
When to See a Doctor
Seek medical attention if:
  • Twitching persists for more than a few weeks.
  • It affects other parts of the face.
  • There is difficulty opening the eye.
  • You experience vision changes or pain.
Conclusion
While most cases of eye twitching are harmless and temporary, addressing lifestyle factors such as sleep, stress, and nutrition can help alleviate symptoms. If twitching persists despite treatment, medical interventions—including Botox or, in rare cases, surgery—may be necessary. Consulting a healthcare provider is essential for persistent or severe cases to rule out underlying conditions.
Appendix D Use of Large Language Models in Paper Writing
We used LLMs to assist with language polishing and minor formatting of the paper. No LLMs were involved in research ideation.
Appendix E Additional Analyses and Clarifications
E.1 Effect of Prompt Difficulty
PiKa adopts a difficulty-driven construction strategy. To examine whether prompt difficulty causally improves alignment performance, we apply a down-quality reconstruction procedure to PiKa prompts and regenerate lower-difficulty variants using the guided template shown below.
System Prompt: You are an expert evaluator for assessing the difficulty level of instructions/questions. Please evaluate the given instruction on a scale of 1–10 based on: Complexity of the task or question Level of domain knowledge required Cognitive load needed to understand and process Technical depth and sophistication Scoring guidelines: 1–2: Very basic, elementary level knowledge 3–4: Intermediate, requires some specialized knowledge 5–6: Advanced, requires solid domain expertise 7–8: Expert level, requires deep specialized knowledge 9–10: Cutting-edge, requires highly specialized or research-level expertise Now, instead of giving a score, please rewrite the given instruction into a new version that would have a difficulty score of 1–2 according to the above scale. Guidelines: Keep the topic and domain consistent Simplify the reasoning and remove complexity Avoid multi-step or abstract reasoning Output strictly in JSON: { ”instruction”: ”¡rewritten easy instruction¿” }
As shown in Table 3, PiKa exhibits substantially higher prompt difficulty than Magpie (average 7.39 vs. 2.65). To more directly test the causal effect of difficulty, we reconstruct lower-quality prompts and regenerate responses using the above template. Each level reduces prompt difficulty while holding other factors fixed. The “mid” variant is derived from the vanilla prompt, and the “low” variant is derived from the “mid” prompt. Difficulty is re-evaluated with GPT-4o.
$\begin{split}P^{\prime}_{\text{new}}&=\mathrm{LLM}(P_{\text{vanilla}},\text{score\_guide}),\\ R_{\text{new}}&=\mathrm{LLM}(P^{\prime}_{\text{new}}).\end{split}$ (7)
The results in Table 3 confirm that higher prompt difficulty leads to stronger alignment gradients.
Magpie-Pro Instruction: ”Write a poem about the beauty of the ocean.” Difficulty Score: 2
PiKa Instruction: ”Analyze the impact of international sanctions and oil price fluctuations on the Russian economy in 2015. Discuss how these factors influenced inflation, recession risks, and consumer prices, particularly in relation to the agricultural and food import ban. Additionally, evaluate the potential for massive withdrawals from the banking system and their broader economic consequences.” Difficulty Score: 7
Figure 6: Example instructions from Magpie-Pro and PiKa datasets. PiKa instructions are generally more complex and domain-specific, while Magpie-Pro tends to focus on simpler, open-ended prompts. We also include detailed instruction-response pairs in Appendix C.
E.2 Effect of Persona Conditioning
Persona conditioning encourages generation of more specialized, higher-difficulty, and less redundant prompts. Removing persona information substantially reduces both difficulty and diversity. Results are shown in Table 6.
Table 6: Persona ablation (10k samples). Mean MND (Minimum Neighbor Distance) measures instruction redundancy (higher is better).
Model (10k) Persona Difficulty Mean MND LC (%) WR (%)
LLaMA-3-8B-PiKa none 3.11 0.0234 13.84 15.53
LLaMA-3-8B-PiKa enabled 7.39 0.497 31.01 30.32
E.3 Persona Domain Distribution
We analyze persona-domain coverage across the PiKa-30k dataset. Table 7 shows that personas largely fall within broad, widely applicable domains rather than narrow or niche specialties.
Table 7: Top 10 persona-domain distribution in PiKa (percentage).
Domain Percentage
General 36.3%
Medicine 16.0%
Biology 13.3%
History 12.5%
Engineering 6.8%
Environmental Science 6.5%
Education 5.8%
Psychology 4.3%
Computer Science 4.0%
To assess robustness to domain composition, we train three PiKa-10k subsets sampled with different seeds. Variance across runs is minimal.
Table 8: Robustness across different random samples of PiKa (10k).
Experiment LC (%) WR (%)
1 31.01 30.32
2 30.09 30.17
3 30.45 30.06
E.4 Evaluation Bias Across Different Judges
To examine whether using GPT-4o as an evaluator introduces systematic bias, we further evaluate PiKa, LLaMA-3-Instruct, and Magpie-Pro using two independent judges: GPT-5 and GPT-4.1-mini. Results in Table 9 show that PiKa consistently outperforms baselines across all evaluators, indicating robustness to choice of judge.
Table 9: Evaluation across different judges.
Judge Model LC (%) WR (%) SD N Avg Len
GPT-5 LLaMA-3-8B-PiKa 19.28 19.00 1.35 805 1927
LLaMA-3-8B-Instruct 15.95 16.56 1.28 805 1949
LLaMA-3-8B-Magpie-Pro 9.51 14.55 1.19 805 2461
GPT-4.1-mini LLaMA-3-8B-PiKa 29.02 27.55 1.52 805 1927
LLaMA-3-8B-Instruct 25.44 25.53 1.45 805 1949
LLaMA-3-8B-Magpie-Pro 20.17 28.54 1.53 805 2461
E.5 Length Analysis Across 5,000 Prompts
We analyze 5,000 prompts, each paired with five candidate responses, to examine whether verbosity bias affects reward-model preferences. Reported lengths correspond to character lengths (approximately 1,300 tokens). Across all prompts, the chosen responses are on average only 7.5% longer than the rejected ones (5,343 vs. 4,969 characters). Moreover, only 30.8% of prompts select the longest candidate as the preferred response, and the global correlation between response length and reward score remains mild ($r=0.31$). These observations collectively indicate that length is not the dominant factor driving reward-model decisions.
Table 10: Length analysis across 5,000 prompts.
Metric Chosen Response Rejected Response $\Delta\%$ Longest Selected (%) Corr (r)
Length (chars) 5,343 4,969 +7.5% 30.8% 0.31
E.6 Robustness Across Different Reward Models
To examine whether PiKa’s evaluation results depend on a particular reward model, we additionally compare two open-source reward models in the 10K setting: Skywork-Reward-V2-LLaMA-3.1-8B and ArmoRM-Llama3-8B-v0.1. As shown in Table 11, the resulting AlpacaEval 2 scores are highly consistent across the two choices. Skywork-Reward-V2-LLaMA-3.1-8B achieves 31.01 LC and 30.32 WR, while ArmoRM-Llama3-8B-v0.1 obtains 29.85 LC and 30.05 WR. The small performance gap suggests that PiKa’s gains are not tied to a single evaluator and remain stable across different open-source reward models.
Reward Model LC WR
Skywork-Reward-V2-LLaMA-3.1-8B 31.01 30.32
RLHFlow/ArmoRM-Llama3-8B-v0.1 29.85 30.05
Table 11: AlpacaEval 2 results under different open-source reward models in the 10K setting.
E.7 Robustness Across Multiple LLM Judges
To assess whether the data-quality conclusions are sensitive to the choice of judge, we further evaluate PiKa and Magpie-Pro using multiple LLM judges, including GPT-4o, Gemini-2.5-Flash, and Gemini-3-Flash. As shown in Table 12, the same overall trend holds across all three judges: PiKa is consistently rated as substantially more difficult than Magpie-Pro, while also achieving slightly higher quality scores on average. These results suggest that the data-quality comparison remains robust across different judges.
Dataset GPT-4o Gemini-2.5-Flash Gemini-3-Flash Avg.
Difficulty score
PiKa 7.39 7.98 6.45 7.27
Magpie-Pro 2.65 4.51 3.17 3.44
Quality score
PiKa 9.57 9.53 9.19 9.43
Magpie-Pro 9.01 9.37 9.03 9.14
Table 12: Data-quality evaluation results across multiple LLM judges. PiKa is consistently rated as more difficult than Magpie-Pro, while also achieving slightly higher quality scores on average.
E.8 Comparison with Difficulty-Matched Filtering
A natural question is whether PiKa’s gains can be recovered simply by selecting the highest-scoring subset from an existing synthetic dataset. To examine this, we construct a control set by selecting the top 30K examples from Magpie-Pro using the same Skywork-Reward-V2 reward model. As shown in Table 13, this filtered subset still has a much lower average difficulty than PiKa-30K (3.01 vs. 7.39), indicating that it does not match the higher-difficulty prompt distribution targeted by PiKa. Correspondingly, PiKa-30K still achieves substantially better AlpacaEval 2 LC performance than the filtered Magpie-Pro subset (32.82 vs. 20.86). These results suggest that PiKa’s advantage cannot be explained solely by reward-model filtering over an existing dataset, but instead comes from producing a substantially harder and more effective training distribution.
Dataset Avg. Difficulty LC
Magpie-Pro-Top-30K 3.01 20.86
PiKa-30K 7.39 32.82
Table 13: Comparison between PiKa-30K and a reward-filtered top-30K subset from Magpie-Pro selected by the same Skywork-Reward-V2 model.
E.9 Controlled Comparison with Different Generators
To further isolate the role of the generator, we conduct a controlled comparison under the same expert-level prompt setting by replacing the GPT-4o generator in PiKa with Llama-3-70B-Instruct, which is also used in Magpie. As shown in Table 14, PiKa with Llama-3-70B-Instruct preserves the high prompt difficulty of PiKa, matching PiKa with GPT-4o on difficulty (7.39 vs. 7.39). However, its response quality is substantially lower than PiKa with GPT-4o (7.01 vs. 9.57), and its downstream AlpacaEval 2 performance also drops markedly. In particular, PiKa with Llama-3-70B-Instruct achieves only 14.59 LC and 16.27 WR, compared with 31.01 LC and 30.32 WR for PiKa with GPT-4o. These results suggest that under expert-level, high-difficulty prompts, a weaker generator cannot produce responses of comparable quality, which further supports the need for a stronger GPT-based generator in our pipeline.
Dataset Difficulty Response Quality LC (%) WR (%)
Magpie-Pro 2.65 9.01 15.42 16.89
PiKa w/ Llama-3-70B-Instruct 7.39 7.01 14.59 16.27
PiKa w/ GPT-4o 7.39 9.57 31.01 30.32
Table 14: Controlled comparison under the same expert-level prompt setting with different generators. Replacing GPT-4o with Llama-3-70B-Instruct preserves prompt difficulty but substantially reduces response quality and downstream chat performance.